40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 14 05:26:39 2012, interferomter tack force, Update, General, daytime tasks 
    Reply  Wed Mar 14 14:09:01 2012, interferomter tack force, Update, General, daytime tasks 
    Reply  Wed Mar 14 16:39:02 2012, Suresh, Update, General, REFL165 signal was not reaching demod board : Fixed Test_CDS_Calibration.png
Message ID: 6418     Entry time: Wed Mar 14 16:39:02 2012     In reply to: 6412
Author: Suresh 
Type: Update 
Category: General 
Subject: REFL165 signal was not reaching demod board : Fixed 

Quote:

The following tasks need to be done in the daytime tomorrow.

  • Hook up the DC output of the Y green BBPD on the PSL table to an ADC channel (Jamie / Steve)
  • Install fancy suspension matrices on PRM and ITMX [#6365] (Jenne)
  • Check if the REFL165 RFPD is healthy or not (Suresh / Koji)
    • According to a simulation the REFL165 demod signal should show similar amount of the signal to that of REFL33.
    • But right now it is showing super tiny signals [#6403]

 The REFL165 RF output was not reaching the Demod board.  The RF cable was disconnected.  I fixed that and then I put in a RF signal at 165MHz , 1.66 mVrms at the test input  (100Hz off set from the 165MHz LO) and saw that the 100 Hz demodulated signal was visible in the dataviewer. 

Test_CDS_Calibration.png

 

Will complete the Optical RF power -> CDS counts calibration tomorrow morning. 

ELOG V3.1.3-